OpenStudy (anonymous):

Solve by the substituion method 3x-7y=43 9x+71=y

OpenStudy (saifoo.khan):

have you tried them? where are u stuck.

OpenStudy (saifoo.khan):

?

hero (hero):

Saifoo, not giving answers anymore?

OpenStudy (saifoo.khan):

Heck no.

hero (hero):

I see it as laziness now.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (chaise):

3x-7y=43 9x+71=y Substitute your y from the second equation into the first equation. 3x-7(9x+71)=43 Now your equation has one variable and can be solved like a normal equation. 3x-7(9x+71)=43 Expand the brackets 3x-63x-497=43 Gather like terms -60x= 540 x = -9 Substitute this back one of the original equations, doesn't matter which one. 9x+71=y 9x-9+71 = y -81+71 = y y = -10